OpenStudy (anonymous):

Evaluating e^(ln√(x))? and lne^5?

OpenStudy (anonymous):

\(e^x\) and \(\ln(x)\) are inverse functions

OpenStudy (anonymous):

in other words \[e^{\ln(\text{whatever})}=\text{whatever}\]

OpenStudy (anonymous):

So the answer for the first is just x?

OpenStudy (anonymous):

similarly \[\ln(e^{\text{anything}})=\text{anything}\]

OpenStudy (anonymous):

Sqrt x

OpenStudy (anonymous):

no, for the first one your "whatever" is \(\sqrt{x}\)

OpenStudy (anonymous):

yeah you got it the second time, that is correct

OpenStudy (anonymous):

Okay, I understand the pattern now. Thank you very much. :)